OpenStudy (anonymous):

Find the x-intercept of the graph of the function show formula work: A(x)=3.2(x-2)+1.8

OpenStudy (lalaly):

x intercept is when A(x)=0 \[3.2x-6.4+1.8=0\]\[3.2x-4.6=0\]\[3.2x=4.6\]\[x=\frac{4.6}{3.2}\]

OpenStudy (lalaly):

x-intercept \[(\frac{4.6}{3.2},0)\]

OpenStudy (anonymous):

Nice skills, I was lost on that one.

OpenStudy (lalaly):

:):)